OpenStudy (anonymous):

Solve the word problem. The base of a triangle with height h and area A is given by the formula b = . What is the base of a triangle with area 17 cm2 and height 5 cm? A. 3.4 cm B. 3.8 cm C. 6.8 cm D.11 cm

OpenStudy (anonymous):

the area of the triangle 1/2 *h*b

OpenStudy (sleepyjess):

Area =\(\dfrac{h*b}{2}\)

OpenStudy (sleepyjess):

so substitute 17 and 5 into the equation and work from there

OpenStudy (anonymous):

b=area*2/h

OpenStudy (anonymous):

b=17*2/5=6.8

OpenStudy (anonymous):

c